# Exact computation of Transfer Entropy with Path Weight Sampling

### v3

@@ Abstract
Jump process. For systems with jumps between a discrete number of states and finite waiting times between the jumps, such as a well-stirred chemical reaction network described by a master equation, or a neural spiking process, the dynamics is governed by a jump propensity matrix $\mathbf{Q}$ of dimensions $\mathcal{N}^{d}\times\mathcal{N}^{d}$, which describes jumps among the $\mathcal{N}$ states of each of the $d$ components of $\mathbf{X}$ 26. In contrast with diffusive processes which must be time-discretized to simulate, jump processes can be simulated exactly with event-driven kinetic Monte-Carlo algorithms such as the Gillespie algorithm27. Hence the transfer of information from $X_{i}$ to $X_{j}$ during the $(k+1)$-th trajectory segment occurs at all instants of time within the segment $X_{j,[k,k+1]}$, rather than at only the endpoint $X_{j}(k+1)$. Therefore in the definition of the stepwise increments to transfer entropy, $\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$, in Eq. 4, the logarithms of the probabilities $P(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]})$ and $P(X_{j}(k+1)|X_{j,[0,k]})$ should be replaced with functionals of the entire $X_{i,[k,k+1]}$ and $X_{j,[k,k+1]}$ segments,
@@
$\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ $=\left\langle\pi_{X_{i}\to X_{j}}-\pi_{X_{j}}\right\rangle$ (11)
where $\alpha$ counts the jumps that change the state of $X_{j}$, $\lambda_{ij}$ and $\lambda_{j}$ are escape propensities for $X_{j}$ in the marginal spaces of $(X_{i},X_{j})$ and $(X_{j})$ respectively, and $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$ similarly are marginal jump propensities for the $\alpha$-th jump2, 15. $Q_{ij},\lambda_{ij}$ and $Q_{j},\lambda_{j}$ are obtained by marginalizing jump propensities from the full $\mathbf{X}$-space into the $(X_{i},X_{j})$ and $(X_{j})$ spaces respectively2, 15 (see SM).
@@
The transfer entropy in Eq. 11 can be calculated with TE-PWS by marginalizing jump propensities in a Monte-Carlo fashion over conditional distributions of hidden variables, similar to the diffusive case. Specifically, the average in Eq. 11 is expressed as a Monte-Carlo average over trajectories of $X_{i}$ and $X_{j}$, and each of $\mathcal{Q}_{ij},\lambda_{ij}$ and $\mathcal{Q}_{j},\lambda_{j}$ is expressed as a Monte-Carlo average over conditional distributions of hidden variables. Hidden variable trajectories are sampled from a $P_{0}$ propagated in the field of the frozen trajectory, and the remaining bias is exactly corrected for through the RR scheme (see SM). Prior work argues that the escape terms involving $\lambda_{ij}$ and $\lambda_{j}$ in Eqs. 9 and 10 respectively can be omitted since they cancel each other on average2. We show however in the SM that the error in the transfer entropy estimate can be reduced by an order of magnitude by exploiting anti-correlated fluctuations between the escape and the jump terms.
Data processing inequality. We demonstrate the utility of TE-PWS by applying it to a three-node motif to study whether transfer entropies obey the Data Processing Inequality (DPI). Unidirectional flow of information between different nodes in a network leads to a DPI for the mutual information28. For a general three-variable process, if the flow of information is $X_{1}\to X_{2}\to X_{3}$, i.e., without feedback, $I(X_{1,[0,N]};X_{3,[0,N]}|X_{2,[0,N]})=0$; here the right arrows denote flow of information as mediated either via activation or repression. This leads by the chain rule to $I(X_{1,[0,N]};X_{3,[0,N]})\leq I(X_{1,[0,N]};X_{2,[0,N]})$28. As transfer entropy equals mutual information in the absence of feedback, it also obeys $\mathcal{T}_{X_{1}\to X_{3}}\leq\mathcal{T}_{X_{1}\to X_{2}}$. This relation bounds the amount of information that can be transmitted from input to output through an intermediate variable, yet is only valid in the absence of feedback. In the presence of an $X_{2}\rightarrow X_{1}$ feedback, i.e., $X_{1}\rightleftarrows X_{2}\to X_{3}$, the mutual information continues to obey its DPI, while the transfer entropy formally doesn’t29, 30. This can be rationalized by considering the limit $X_{1}\leftarrow X_{2}\rightarrow X_{3}$, where $X_{2}$ controls both $X_{1}$ and $X_{3}$. Here the $X_{1}$ trajectories would still be predictive of fluctuations in $X_{3}$, thus $\mathcal{T}_{X_{1}\to X_{3}}\geq 0$ even as $\mathcal{T}_{X_{1}\to X_{2}}=0$. However, where the crossover from the feedforward to the feedback-dominated regime occurs, and to what extent transfer entropies can overcome the DPI in practice, is currently not understood.
Figure 3: (a) and (b) are typical $X_{1}$ (green) and $X_{2}$ (black) trajectories for models A and B at $a_{12}=2a_{21}=0.4$ and $a_{12}=a_{21}=-6$ respectively. (c) Transfer entropy rates in model A, denoted as $\dot{\mathcal{T}}_{X_{1}\to X_{3}}^{(\mathrm{A})}$ and $\dot{\mathcal{T}}_{X_{1}\to X_{2}}^{(\mathrm{A})}$, as a function of an increasing ratio of feedback to feedforward strength $f^{*}=a_{12}/a_{21}$, where the feedforward strength is kept constant at $a_{21}=0.2$. (d) Similar to (c) but for model B, where the feedforward strength is kept constant at $a_{21}=-4$. (e) Ratio of transfer entropy rates overcomes DPI bound (dashed line at $\mathcal{T}^{*}=1$) with increasing feedback.
We implement the motif $X_{1}\rightleftarrows X_{2}\to X_{3}$ in two diffusive models of mutual repression between $X_{1}$ and $X_{2}$, labeled as models A and B. In both models, $X_{3}$ rapidly copies the state of $X_{2}$ such that the information loss from $X_{2}$ to $X_{3}$ is low. Model A is a three-dimensional OU process with linear feedback, where the ratio of the feedback to feedforward spring constants $f^{*}\equiv a_{12}/a_{21}$ is varied to study the violation of DPI. Model B is a nonlinear extension of A inspired by a genetic toggle switch31 where the drifts for $X_{1}$ and $X_{2}$ are changed to $F_{1}=-a_{11}X_{1}-a_{12}(1+X_{1}^{2})/(1+X_{1}^{2}+X_{2}^{2})$ and $F_{2}=-a_{22}X_{2}-a_{21}(1+X_{2}^{2})/(1+X_{1}^{2}+X_{2}^{2})$; we choose $a_{12}/a_{21}=f^{*}$ such that the two models can be compared. Other model parameters are provided in the EM. Typical $X_{1}$ and $X_{2}$ trajectories at high feedback are shown in Figs. 3a and b. Model A only shows regression to the mean for both $X_{1}$ and $X_{2}$, while model B additionally shows switching between (low,high) and (high,low) values of $(X_{1},X_{2})$ in the regime of both strong feedforward and feedback coupling, around $f^{*}=1$.
We plot in Figs. 3c and d transfer entropy rates for the two models. Each data point is obtained from steady state trajectories using TE-PWS. In model A, increasing the feedback strength $f^{*}$ keeps $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ unchanged while $\dot{\mathcal{T}}_{X_{1}\to X_{3}}$ monotonically increases. The former follows analytically from the expression for the transfer entropy in a two-dimensional OU process24 (see EM). The latter is an empirical result that arises because, when $f^{*}$ is increased, $X_{1}$ becomes increasingly correlated with $X_{2}$ while $X_{2}$ continues to be copied accurately by $X_{3}$. In contrast, in model B, both $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ and $\dot{\mathcal{T}}_{X_{1}\to X_{3}}$ peak near the switching regime, phenomenologically similar to how a bistable three-node motif with feedback has been shown to behave through approximate theory15. The values of the transfer entropy rates in model B are also amplified many-fold compared to model A due to stronger correlations between $X_{1}$ and $X_{2}$ resulting from the stronger, nonlinear coupling, with the concomitant switching.
In Fig. 3e we have plotted the amount of DPI violation in both models, quantified as the ratio of transfer entropy rates $\mathcal{T}^{*}\equiv\dot{\mathcal{T}}_{X_{1}\to X_{3}}/\dot{\mathcal{T}}_{X_{1}\to X_{2}}$, as a function of the ratio of the feedback to feedforward strengths, $f^{*}$. Surprisingly, we find that regardless of the nature of the variation of the individual transfer entropy rates with increasing feedback, i.e., monotonic or non-monotonic (Figs. 3c and d), the ratio $\mathcal{T}^{*}$ monotonically increases with increasing feedback (Fig. 3 e). Moreover, the ratio overcomes the DPI bound when the strength of the feedback becomes comparable to that of the feedforward coupling. Our results thus show that when the feedback $X_{2}\to X_{1}$ dominates over the feedforward interaction, $X_{1}\to X_{2}$, the feedforward entropy $\mathcal{T}_{X_{1}\to X_{3}}$ becomes larger than the feedforward entropy $\mathcal{T}_{X_{1}\to X_{2}}$. The mutual information, on the other hand, continues to obey its DPI, $I(X_{1,[0,N]};X_{3,[0,N]})\leq I(X_{1,[0,N]};X_{2,[0,N]})$. We expect further analytical work in the OU process to be able to support this empirical result.
In conclusion, we have developed a method that for the first time makes it possible to compute transfer entropies exactly for any stochastic model. We have shown that TE-PWS is applicable to diverse stochastic processes and can elucidate novel physics of information transmission. We expect transfer entropies computed by TE-PWS to be used as ground truth for a wide range of goals, such as the characterization and design of information flow in natural and engineered information processing systems, and causality detection.
Data availability Python implementations of TE-PWS for diffusive and jump processes, and data that reproduce the findings of this study, are openly available on Zenodo at https://zenodo.org/doi/10.5281/zenodo.1361736532.
Appendix A End Matter
Model details. Figs. 2 a and b show transfer entropies in a minimal stochastic model for gene expression25, 11, 24. The model is a three-dimensional OU process with spring constants $a_{11}=\mu_{E}+\mu_{0}T_{\mu E}(T_{EG}-1)$, $a_{12}=-\mu_{0}(T_{EG}-1)$, $a_{13}=\mu_{0}T_{EG}$, $a_{21}=-T_{\mu E}[\beta_{G}-\mu_{E}-\mu_{0}T_{\mu E}(T_{EG}-1)]$, $a_{22}=-[\mu_{0}T_{\mu E}(T_{EG}-1)-\beta_{G}]$, $a_{23}=-[\beta_{G}-\beta_{\mu}-\mu_{0}T_{\mu E}T_{EG}]$, $a_{31}=a_{32}=0$ and $a_{33}=\beta_{\mu}$, and diffusion constants $D_{11}=D_{E}\mu_{0}^{2}$, $D_{22}=\beta_{\mu}\eta_{\mu}^{2}+\beta_{G}\eta_{G}^{2}+D_{E}\mu_{0}^{2}T_{\mu E}^{2}$, $D_{33}=\beta_{\mu}\eta_{\mu}^{2}$, $D_{12}=D_{21}=D_{E}\mu_{0}^{2}T_{\mu E}$, $D_{13}=D_{31}=0$ and $D_{23}=D_{32}=\beta_{\mu}\eta_{\mu}^{2}$, where the experimentally determined values of the parameters are $\mu_{0}=0.23h^{-1}$, $\beta_{\mu}=\beta_{G}=0.33h^{-1}$, $\beta_{E}=5.63h^{-}1$, $\eta_{E}=1.03$, $\eta_{\mu}=0.16$, $\eta_{G}=0.22$, $T_{EE}=1$, $T_{\mu E}=0.7$, $T_{EG}=1.3$, $\mu_{E}=\mu_{0}(1+T_{\mu E}-T_{EE})$ and $D_{E}=\eta_{E}^{2}/\beta_{E}$.25, 24. We also used $\delta t=9.2\times 10^{-5}\mu_{0}^{-1}$, $M_{1}=96$ and $M_{2}=200$.
For Figs. 2c and d, the model is a two-dimensional OU process with parameters $a_{11}=1$, $a_{22}=2$, $a_{12}=a_{21}=0.5$, $D_{11}=1.5$, $D_{22}=0.5$ and $D_{12}=0$. Additionally, $\delta t=0.01a_{11}^{-1}$, $M_{1}=100$ and $M_{2}=1000$.
For models A and B in Fig. 3, we have chosen $a_{11}=a_{22}=1$, $a_{33}=-a_{32}=2$, $a_{13}=a_{31}=a_{23}=0$ and $D_{ij}=\delta_{ij}$. Additionally, $\delta t=0.01a_{11}^{-1}$, $M_{1}=100$ and $M_{2}=1000$.
In all simulations in the main text, trajectories have been initialized from a histogram of the corresponding steady-state distributions. Error bars for all results are computed as twice the standard deviations using multiple sets of statistically independent simulations.
where $r_{2}=[a_{11}^{2}+(D_{11}/D_{22})a_{21}^{2}-2(D_{12}/D_{22})a_{11}a_{21}]^{1/2}$. For model A in the section Data processing inequality of the main text, the dynamics of $X_{1}$ and $X_{2}$ comprises that of a two-dimensional OU process because there is no feedback from $X_{3}$. Hence the transfer entropy rate $\dot{\mathcal{T}}_{X_{1}\to X_{2}}$ is given by Eq. 12, which demonstrably does not depend on the $X_{2}\to X_{1}$ feedback strength $a_{12}$, i.e., on $f^{*}$ (see Fig. 3c).
@@ Appendix A End Matter
where $|\mathbf{D}|$ is the determinant of the diffusion constant matrix. This form also holds for systems with inertia if the generalized coordinate vector $\mathbf{X}$ contains both positions and velocities33, 34.
Jump processes can be simulated with a Gillespie algorithm, which is exact, i.e., does not make a timestep error27. To clarify the quantities in the main text (Eqs. 9-11), and elucidate how they can be obtained in the Gillespie simulations, we briefly describe the scheme. At any given time, the waiting time till the next jump is exponentially distributed. Hence at the start of the trajectory and after every jump, denoting the time as $t$, we sample a uniform random number $u\in[0,1]$ and determine the wait time $\tau$ by solving
@@
for $\tau$. Here $\lambda(t^{{}^{\prime}})$ is the escape propensity from state $\mathbf{X}(t^{{}^{\prime}})$, given by the sum of the jump propensities $\mathcal{Q}_{\beta}(t^{{}^{\prime}})$ taking the system out of the state $\mathbf{X}(t^{{}^{\prime}})$ at time $t^{{}^{\prime}}$. When propagating all coordinates at once, $\lambda(t^{{}^{\prime}})$ is independent of $t^{{}^{\prime}}$ in between jumps; hence Eq. 14 is trivially solved for $\tau$. Then a second uniform random number is drawn to choose which jump to fire, with probabilities proportional to their jump propensities. The probability density of a trajectory segment $\mathbf{X}_{[t,t+\delta t]}$ is written as
[…]
where $\alpha$ sums over all jumps in the full trajectory, $N_{\mathrm{tot}}$ in number, occurring at times $t_{\alpha}$.
@@
    For each pair of $(X_{i}^{(\nu)},X_{j}^{(\nu)})$ trajectories, propagate $M_{2}$ trajectories of hidden variables $X_{l}$ using the chosen reference distribution $P_{0}(X_{l,[0,k]})$. Initial conditions should be sampled from the same joint distribution $P(X_{i}^{(\nu)}(0),X_{j}^{(\nu)}(0),X_{l}(0))$ that $X_{i}^{(\nu)}(0)$ and $X_{j}^{(\nu)}(0)$ were drawn from.
@@
    After every $\delta t$ time, recalculate logarithmic weights $w^{(\mu)}$ for the trajectories defined as the logarithm of the ratio between the joint distribution $P(X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]})$ and the reference distribution $P_{0}(X_{l,[0,k]})$, using Eqs. A and 14. This ratio is proportional to the ratio between the conditional distribution $P(X_{l,[0,k]}|X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)})$ and the reference distribution $P_{0}(X_{l,[0,k]})$, as proven in the SM. Then calculate the uniformity in the weights with a uniformity parameter $\kappa=(\sum_{\mu}\exp w^{(\mu)})^{2}/\sum_{\mu}\exp(2w^{(\mu)})$, where sums of exponentials of weights are always performed with the Log-Sum-Exp trick35.
@@
    Calculate the contribution to the transfer entropy in the space of $(X_{i},X_{j})$ at the $(k+1)$-th step, denoted as $\mathcal{T}_{a}^{(\nu)}[k]$. This is done by computing the average in Eq. 8 for diffusive processes, and taking an expectation of Eq. 9 for jump processes (see SM).
@@
    Akin to steps 2-5, for each $(X_{j}^{(\nu)})$ trajectory, propagate $M_{2}$ trajectories of $(X_{i},X_{l})$ using reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$; accumulate weights for each trajectory as the ratio between the joint distribution $P(X_{i,[0,k]},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]})$ and the reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$; compute the contribution to the transfer entropy in the $(X_{j})$ space at the $(k+1)$-th step, denoted as $\mathcal{T}_{b}^{(\nu)}[k]$ (Eqs. 7 and 10); and resample if the uniformity in the weights is low, i.e., $\kappa<M_{2}/2$.
@@
    Finally, compute $\mathcal{T}_{X_{i}\to X_{j}}$ by combining $\mathcal{T}_{a}^{(\nu)}$ and $\mathcal{T}_{b}^{(\nu)}$ from all timesteps using Eqs. 5 and 11 for diffusive and jump processes respectively.
For clarity, a pseudocode to compute transfer entropy $\mathcal{T}_{X_{1}\to X_{3}}$ in a three-variable process is given in the SM. We note that TE-PWS is easily parallelized as it is a Monte-Carlo algorithm. Specifically, steps 1-7 can be executed independently for each of the $M_{1}$ trajectories mentioned in step 1, using parallel processors. The transfer entropy is computed at the end as an average over the results from all processors.
@@
November 25, 2024
Appendix B Pseudocode

### v4

@@ Appendix A End Matter
Jump process. For systems with jumps between a discrete number of states and finite waiting times between the jumps, such as a well-stirred chemical reaction network described by a master equation, or a neural spiking process, the dynamics is governed by a jump propensity matrix $\mathbf{Q}$ of dimensions $\mathcal{N}^{d}\times\mathcal{N}^{d}$, which describes jumps among the $\mathcal{N}$ states of each of the $d$ components of $\mathbf{X}$35. In contrast with diffusive processes which must be time-discretized to simulate, jump processes can be simulated exactly with event-driven kinetic Monte-Carlo algorithms such as the Gillespie algorithm36. Hence the transfer of information from $X_{i}$ to $X_{j}$ during the $(k+1)$-th trajectory segment occurs at all instants of time within the segment $X_{j,[k,k+1]}$, rather than at only the endpoint $X_{j}(k+1)$. Therefore in the definition of the stepwise increments to transfer entropy, $\mathcal{T}^{[k]}_{X_{i}\to X_{j}}$, in Eq. 5, the logarithms of the probabilities $P(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]})$ and $P(X_{j}(k+1)|X_{j,[0,k]})$ should be replaced with functionals of the entire $X_{i,[k,k+1]}$ and $X_{j,[k,k+1]}$ segments,
@@
$\mathcal{T}^{[k]}_{X_{i}\to X_{j}}$ $=\left\langle\pi_{X_{i}\to X_{j}}-\pi_{X_{j}}\right\rangle$ (14)
@@
where $\alpha$ counts the jumps that change the state of $X_{j}$, $\lambda_{ij}$ and $\lambda_{j}$ are escape propensities for $X_{j}$ in the marginal spaces of $(X_{i},X_{j})$ and $(X_{j})$, respectively, and $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$ similarly are marginal jump propensities for the $\alpha$-th jump2, 15. $Q_{ij},\lambda_{ij}$ and $Q_{j},\lambda_{j}$ are obtained by marginalizing jump propensities from the full $\mathbf{X}$-space into the $(X_{i},X_{j})$ and $(X_{j})$ spaces, respectively2, 15. The marginalization can be performed with TE-PWS in a Monte-Carlo fashion over conditional distributions of hidden variables, similar to the diffusive case (see SM-E). Prior work argues that the escape terms involving $\lambda_{ij}$ and $\lambda_{j}$ in Eqs. 12 and 13, respectively, can be omitted since they cancel each other on average2. We show however in SM-F that the error in the transfer entropy estimate can be reduced by an order of magnitude by exploiting anti-correlated fluctuations between the escape and the jump terms.
Table 1: Systematic errors and computational costs of approximate methods in the linear model A and the nonlinear model D. The input data size has been kept comparable whenever possible within non-prohibitive computational costs.
Exact Gaussian KSG $(k\to\infty)$ KSG $(k=0)$
Linear model Input data size $NM_{1}$ $1.6\times 10^{6}$ $1.6\times 10^{6}$ $1.6\times 10^{5}$ $1.6\times 10^{6}$
(model A) $\dot{\mathcal{T}}_{1\to 2}~{}[10^{-2}~{}\mathrm{nats~{}}a_{11}]$ $3.98\pm 0.06$ $3.90\pm 0.33$ $5.82\pm 0.07$ $8.98\pm 3.39$
CPU hours $\mathcal{C}$ $1.88$ $1.35$ $75.92$ $9.55$
Nonlinear model Input data size $NM_{1}$ $1.6\times 10^{6}$ $1.6\times 10^{6}$ $1.28\times 10^{5}$ $1.6\times 10^{6}$
(model D) $\dot{\mathcal{T}}_{1\to 3}~{}[10^{-2}~{}\mathrm{nats~{}}a_{11}]$ $4.23\pm 0.10$ $5.33\pm 0.46$ $2.25\pm 0.04$ $2.61\pm 3.48$
CPU hours $\mathcal{C}$ $1.90$ $19.00$ $71.50$ $9.13$
@@
where $|\mathbf{D}|$ is the determinant of the diffusion constant matrix. This form also holds for systems with inertia if the generalized coordinate vector $\mathbf{X}$ contains both positions and velocities37, 38.
Jump processes can be simulated with a Gillespie algorithm, which is exact, i.e., does not make a timestep error36. The probability density of a trajectory segment $\mathbf{X}_{[t,t+\delta t]}$ is written as
@@
where $\lambda(t^{{}^{\prime}})$ is the escape propensity from state $\mathbf{X}(t^{{}^{\prime}})$, given by the sum of the jump propensities $\mathcal{Q}_{\beta}(t^{{}^{\prime}})$ taking the system out of the state $\mathbf{X}(t^{{}^{\prime}})$ at time $t^{{}^{\prime}}$, and $\alpha$ sums over all jumps in the full trajectory, $N_{\mathrm{tot}}$ in number, occurring at times $t_{\alpha}$.
@@
    For each pair of $(X_{i}^{(\nu)},X_{j}^{(\nu)})$ trajectories, propagate $M_{2}$ trajectories of hidden variables $X_{l}$ using the chosen reference distribution $P_{0}(X_{l,[0,k]})$. Initial conditions should be sampled from the same joint distribution $P(X_{i}(0),X_{j}(0),X_{l}(0))$ that $X_{i}^{(\nu)}(0)$ and $X_{j}^{(\nu)}(0)$ were drawn from.
@@
    After every $\delta t$ time, recalculate logarithmic weights $w^{(\mu)}$ for the trajectories defined as the logarithm of the ratio between the joint distribution $P(X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]})$ and the reference distribution $P_{0}(X_{l,[0,k]})$, using Eqs. A and 16. This ratio is proportional to the ratio between the conditional distribution $P(X_{l,[0,k]}|X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)})$ and the reference distribution $P_{0}(X_{l,[0,k]})$, as proven in SM-C. Then calculate the uniformity in the weights with a uniformity parameter $\kappa=(\sum_{\mu}\exp w^{(\mu)})^{2}/\sum_{\mu}\exp(2w^{(\mu)})$, where sums of exponentials of weights are always performed with the Log-Sum-Exp trick39.
@@
    Calculate the contribution to the transfer entropy in the space of $(X_{i},X_{j})$ at the $(k+1)$-th step, denoted as $\mathcal{T}_{a}^{(\nu)}[k]$. This is done by computing the average in Eq. 9 for diffusive processes, and taking an expectation of Eq. 12 for jump processes (see SM-E).
@@
    Akin to steps 2-5, for each $(X_{j}^{(\nu)})$ trajectory, propagate $M_{2}$ trajectories of $(X_{i},X_{l})$ using reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$; accumulate weights for each trajectory as the ratio between the joint distribution $P(X_{i,[0,k]},X_{j,[0,k]}^{(\nu)},X_{l,[0,k]})$ and the reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$; compute the contribution to the transfer entropy in the $(X_{j})$ space at the $(k+1)$-th step, denoted as $\mathcal{T}_{b}^{(\nu)}[k]$ (Eqs. 8 and 13); and resample if the uniformity in the weights is low, i.e., $\kappa<M_{2}/2$.
@@
    Finally, compute $\mathcal{T}_{X_{i}\to X_{j}}$ by combining $\mathcal{T}_{a}^{(\nu)}$ and $\mathcal{T}_{b}^{(\nu)}$ from all timesteps using Eqs. 6 and 14 for diffusive and jump processes, respectively.
For clarity, a pseudocode to compute the transfer entropy $\mathcal{T}_{X_{1}\to X_{3}}$ in a three-variable process is given in SM-B. We note that TE-PWS is easily parallelized as it is a Monte-Carlo algorithm. Specifically, steps 1-7 can be executed independently for each of the $M_{1}$ trajectories mentioned in step 1, using parallel processors. The transfer entropy is computed at the end as an average over the results from all processors.
@@
July 10, 2025
Appendix B SM-A: Simulation details
Model A. For Fig. 2a in the main text, the model (model A) is a two-dimensional OU process with parameters $a_{11}=a_{22}=1$, $a_{12}=a_{21}=0.9$, $D_{11}=0.2$, $D_{22}=1$ and $D_{12}=0$. Additionally, $\delta t=0.01a_{11}^{-1}$, $M_{1}=10^{6}$ and $M_{2}=100$ for TE-PWS. For the Gaussian framework, our protocol for obtaining the converged transfer entropy rates in the infinite data limit are described in the section SM-H: Comparison with approximate methods.
Model B. Figs. 2b and c in the main text show transfer entropies in the minimal stochastic model for gene expression from Refs. 1, 2, 3. The model (model B) is a three-dimensional OU process with spring constants $a_{11}=\mu_{E}+\mu_{0}T_{\mu E}(T_{EG}-1)$, $a_{12}=-\mu_{0}(T_{EG}-1)$, $a_{13}=\mu_{0}T_{EG}$, $a_{21}=-T_{\mu E}[\beta_{G}-\mu_{E}-\mu_{0}T_{\mu E}(T_{EG}-1)]$, $a_{22}=-[\mu_{0}T_{\mu E}(T_{EG}-1)-\beta_{G}]$, $a_{23}=-[\beta_{G}-\beta_{\mu}-\mu_{0}T_{\mu E}T_{EG}]$, $a_{31}=a_{32}=0$ and $a_{33}=\beta_{\mu}$, and diffusion constants $D_{11}=D_{E}\mu_{0}^{2}$, $D_{22}=\beta_{\mu}\eta_{\mu}^{2}+\beta_{G}\eta_{G}^{2}+D_{E}\mu_{0}^{2}T_{\mu E}^{2}$, $D_{33}=\beta_{\mu}\eta_{\mu}^{2}$, $D_{12}=D_{21}=D_{E}\mu_{0}^{2}T_{\mu E}$, $D_{13}=D_{31}=0$ and $D_{23}=D_{32}=\beta_{\mu}\eta_{\mu}^{2}$, where the experimentally determined values of the parameters are $\mu_{0}=0.23h^{-1}$, $\beta_{\mu}=\beta_{G}=0.33h^{-1}$, $\beta_{E}=5.63h^{-}1$, $\eta_{E}=1.03$, $\eta_{\mu}=0.16$, $\eta_{G}=0.22$, $T_{EE}=1$, $T_{\mu E}=0.7$, $T_{EG}=1.3$, $\mu_{E}=\mu_{0}(1+T_{\mu E}-T_{EE})$ and $D_{E}=\eta_{E}^{2}/\beta_{E}$.1, 3. We also used $\delta t=9.2\times 10^{-5}\mu_{0}^{-1}$, $M_{1}=96$ and $M_{2}=200$.
Models C and D. For models C and D in Fig. 3 in the main text, we have chosen $a_{11}=a_{22}=1$, $a_{33}=-a_{32}=2$, $a_{13}=a_{31}=a_{23}=0$ and $D_{ij}=\delta_{ij}$. Additionally, $\delta t=0.01a_{11}^{-1}$, $M_{1}=100$ and $M_{2}=1000$.
In all simulations, trajectories have been initialized from a collection of the corresponding steady-state distributions, including for marginalization in TE-PWS. For details of initializing TE-PWS trajectories, see the SM section SM-D: Choice of reference distribution. Error bars for all results are computed as twice the standard deviations using multiple sets of statistically independent simulations. For comparison with the performance of TE-PWS, the KSG method with 4 nearest neighbors (the ‘first’ KSG algorithm in [4]) was implemented with the openly available JIDT toolkit5. The computational cost of all techniques was monitored, for a total of 40 independent samples for each technique, with the openly available process analysis workbench Procpath6. The CPU cost mentioned in Table 1 of the main text corresponds to a calculation for only the largest studied history length $(k+1)$ in each case. For details on the convergence of the transfer entropy rate from Gaussian and the KSG algorithms (data presented in Table 1 of the main text) with respect to the history length $(k+1)$ and the data size $NM_{1}$, see the SM section SM-H: Comparison with approximate methods.
Appendix C SM-B: Pseudocode